A uniform 80.0 N ladder 4.0 m long is placed against a frictionless wall with its base situated 2.0 from the wall. Find the forces exerted by the wall and by the ground on the ladder.


1
Expert's answer
2021-02-19T09:43:43-0500

Let normal reaction at endpoints be "N_1, N_2"

Force of friction acting between ground and wall is f.

The weight of the ladder will act at the mid-point of the ladder.

Let angle made by the wall with ground is "\theta, cos\theta = \frac{2}{4}= \frac{1}{2}"


Then for equilibrium in the horizontal and vertical direction

"f = N_2" (1)

"N_1 = mg = 80N" (2)


Taking torque about point on ground,

"N_2 lsin\theta = mg\frac{l}{2} cos\theta \implies N_2 = 23.1N"